OpenStudy (anonymous):

the force of gravitation between two objects is 500 newtons. a) what would the force of gravitation be if the distance between these two objects increased by times? b)suppose the distance stayed the same as it was origmally but the mass of one of the bodies doubled. what would the new force of gravitational attraction be?

OpenStudy (anonymous):

\[F= \frac{gM_1M_2}{r^2}\]

OpenStudy (anonymous):

assume that in your case if r=1, then F=500 so if r = x, F=?

OpenStudy (anonymous):

500

OpenStudy (anonymous):

no idea, you left out the multiplier

OpenStudy (anonymous):

basically assume the gM1M2 = 500 so just plug in the necessary multiplier of r

OpenStudy (anonymous):

5n would be the answer?

OpenStudy (anonymous):

I DONT KNOW YOU LEFT OUT A NUMBER reread your problem please

OpenStudy (anonymous):

500/(10)2=500/100=5n